OpenStudy (lgg23):

For the following series, does it diverge or converge? If it converges, what is the limit? Sum from n=1 to inf of ln(2(n+1))-ln(2n)

OpenStudy (anonymous):

This seems to diverge. I would break up into two sums:\[\sum_{1}^{\infty} \ln(2(n+1)) - \sum_{1}^{\infty} \ln(2n)\]Remember that if we have two sums, if one diverges, the sum diverges. So, we have to prove that either sum diverge (both diverge, pick one and solve it).

OpenStudy (anonymous):

Typo: I meant, if we have a sum of two sums, if one of the sums diverges, the sum of sums diverges, i.e.,\[\sum_{i = 1}^{\infty} a_{i} + \sum_{i=1}^{\infty} b_i\]if {a_i} or {b_i} diverges, their sum diverge.

OpenStudy (lgg23):

Thank you.